OpenStudy (anonymous):

Can anybody help me verifying this identity: cot(x-π/2) = -tan x

OpenStudy (anonymous):

\[\begin{align*} \cot\left(x-\frac{\pi}{2}\right)&=\frac{\cos\left(x-\dfrac{\pi}{2}\right)}{\sin\left(x-\dfrac{\pi}{2}\right)}\\\\ &=\frac{\cos x\cos\dfrac{\pi}{2}+\sin x\sin\dfrac{\pi}{2}}{\sin x\cos\dfrac{\pi}{2}-\cos x\sin\dfrac{\pi}{2}}\\\\ &=\frac{\sin x}{-\cos x}\\\\ &=-\tan x \end{align*}\]